# Role of sub-barrier dynamics in coherent and incoherent radiation during tunneling ionization

### v1

@@ Abstract
Radiation during strong-field tunneling ionization is investigated. The spontaneous as well as the coherent component of the radiation are calculated and the role of the under-the-barrier dynamics is analyzed. The competition of the contribution of different mechanisms in different spectral regions is examined. Transition-like radiation (Brunel radiation) is dominated at low-frequencies, Thomson scattering – at the laser frequencies, and radiative recombination via the three-step process – at high-order harmonics. The under-the-barrier dynamics is shown to enhance the Brunel radiation for near-zero-frequencies in asymmetric laser pulses for spontaneous as well as for the coherent radiation.
@@ I Introduction
It is well known that tunneling ionization is accompanied by high-order harmonic generation (HHG) via the three-step model [1], when the recombination of the ionized electron wave packet to the bound state at the recollision after the excursion in the laser field and gaining energy results in the emission of a high energy photon [2, 3, 4]. At the recollision, radiation may emerge also due to electron scattering, rather than recombination, yielding free-free HHG [5]. In the macroscopic picture, the ionization of a gas target results in the creation of an electron current after the interaction with a laser pulse, which can be a source of low frequency THz radiation, with the emitted frequency determined by the laser pulse envelope [6, 7, 8, 9]. The same mechanism can work also in the case of a single atom producing THz radiation in asymmetric laser pulses, e.g. in a two-color laser field [10, 11].
@@
In contrast to the coherent radiation discussed above, emission of spontaneous low-order harmonic radiation during the bound-continuum transition with the origin similar to Brunel radiation [12, 13] has been recently shown in Ref. [14], with the typical frequency determined by the Keldysh time [15]. Finally, harmonic radiation at tunneling ionization arises also from Thomson scattering of the laser wave off the ionized free electron wave packet, with higher harmonics than the fundamental one suppressed in the nonrelativistic regime.
@@
Coherent radiation should be distinguished from the incoherent one of the same origin. For instance, the spontaneous recombination of a high-energy electron to an atomic bound state results in a photon emission of a random phase, which is an incoherent process [16]. However, when the recombining electron is released from the bound state and recombines into the same bound state driven by a laser field, the HHG is coherent with the phase determined by the coherent superposition of the ground state wave function with the ionization continuum wave packet. Similarly, the coherent part of radiation due to Compton scattering of a strong laser wave from an ionized electron wave packet [17, 18] is phase-matched in forward direction, while the incoherent one has a broad angular distribution [19, 20, 21, 22, 23]. For the incoherent process, in contrast to the coherent one, the final state of the electron does not exist initially. The coherent radiation intensity scales as $\sim{\cal N}^{2}$ with the number of emitters $\cal N$}, due to the phase matching of the radiation, while for the incoherent radiation interference is suppressed and the total intensity scaling is $\sim{\cal N}$ [24]. The coherent radiation can be calculated using the quantum mechanical expectation value of the current density within the classical radiation theory [19].
@@
In this paper we investigate the role of the under-the-barrier dynamics for the spontaneous and coherent Brunel radiation during tunneling ionization. We analyze how the different mechanisms of the radiation during the strong field ionization are scaled in different spectral regions of emission and the feasibility of their observation in a coherent or incoherent form.
@@
The structure of the paper is the following. In Sec. II the coherent radiation during tunneling ionization is calculated using the strong field approximation (SFA). Spontaneous emission at the tunneling ionization is discussed in Sec. III. We give estimations of radiation intensities in Sec. IV. Our conclusion is given in Sec. V.
@@ II Coherent radiation during tunneling ionization
Figure 1: Coherent radiation spectra during tunneling ionization with $E_{0}=0.1$, $\omega_{0}=0.05$, and $N=6$: (a) in the asymmetric laser pulse, (b) in the symmetric laser pulse. Comparison of the Brunel radiation via $j_{11}$ with the three-step HHG ($j_{01}+j_{10}$) and with the Babushkin model ($j_{B}$). The total radiation with the full current $j=j_{01}+j_{10}+j_{11}$ is given for comparison.
@@
We calculate coherent radiation emitted during tunneling ionization of a laser-driven hydrogenlike atom for the nonrelativistic regime. The dynamics is described by the Schrödinger equation $i\partial_{t}=H$ with Hamiltonian
$H=\mathbf{p}^{2}/2+V(r)+\mathbf{r}\cdot\mathbf{E}(t)+\mathbf{r}\cdot\hat{\mathbf{E}}(t),$ (1)
where $V(\mathbf{r})=2\pi/\kappa\delta(\mathbf{r})(1+\partial_{r})$ is the regularized atomic zero-range potential, $\mathbf{E}(t)$ the laser electric field, $\hat{\mathbf{E}}(\mathbf{r},t)=\sum_{\mathbf{k}}\sqrt{\frac{2\pi\omega}{\cal{V}}}\hat{\mathbf{e}}\left(a_{\mathbf{k}}e^{i\mathbf{k}\cdot\mathbf{r}-i\omega t}+a^{+}_{\mathbf{k}}e^{i\mathbf{k}\cdot\mathbf{r}-i\omega t}\right)$ is the field of the emitted photons in second quantization, with the creation and annihilation operators $a^{+}_{\mathbf{k}}$ and $a_{\mathbf{k}}$, the emitted photon frequency $\omega=c|\mathbf{k}|$, the polarization unit vector $\hat{\mathbf{e}}$, the speed of light $c$, and the quantization volume $\cal{V}$. Atomic units are used throughout. In the nonrelativistic regime the dipole approximation for the emission is applied, neglecting the spatial dependence of the emission field $\mathbf{k}\cdot\mathbf{r}\ll 1$.
With the electron wave function $\Psi$ as a solution of the Schrödinger equation in the atomic potential and the laser electric field, the amplitude of the coherent photon emission process in the first-order perturbation will read
$m_{\mathbf{k}}=-i\int ds\langle\Psi(s),1_{\mathbf{k}}|\mathbf{r}\cdot\hat{\mathbf{E}}(s)|\Psi(s),0_{\mathbf{k}}\rangle,$ (2)
yielding the differential probability for the coherent photon emission:
$\frac{dw}{d^{3}\mathbf{k}}=\left|m_{\mathbf{k}}\right|^{2}\frac{\cal{V}}{(2\pi)^{3}}.$ (3)
Describing the interaction of the electron with the photon field in the velocity gauge:
$m_{\mathbf{k}}=-i\int ds\langle\Psi(s),1_{\mathbf{k}}|\hat{\mathbf{A}}(s)\cdot\hat{\mathbf{p}}|\Psi(s),0_{\mathbf{k}}\rangle,$ (4)
where $\hat{\mathbf{A}}=-\hat{\mathbf{E}}(t)/\omega$ is the vector potential operator for the emitted photon.
After projecting on the photon initial and final states, and integrating over the emission direction, we obtain
$\frac{dw}{d\omega}=\frac{2\alpha\omega}{3\pi c^{2}}\left|\int ds\langle\Psi(s)\left|\hat{\mathbf{p}}\right|\Psi(s)\rangle\exp(i\omega s)\right|^{2}$ (5)
with the fine-structure constant $\alpha$. The spectral intensity of Eq. (5) can be viewed as classical radiation of the quantum current $\mathbf{j}(\mathbf{r},t)=\Psi(\mathbf{r},t)^{*}\hat{\mathbf{p}}\Psi(\mathbf{r},t)$. In this way the coherent part of the radiation is derived [24].
Figure 2: Coherent radiation spectra during tunneling ionization in the asymmetric laser pulse: (a,c) $E_{0}=0.1$ [zoom in Fig. 1(a)]; (b,d) $E_{0}=0.075$. Comparison of the Brunel radiation via $j_{11}$ with the three-step HHG ($j_{01}+j_{10}$) and with the Babushkin model ($j_{B}$). (c,d) Zoom in the low energy spectral region $\omega<0.2$ eV.
Figure 3: The role of the under-the-barrier dynamics for coherent radiation in a half-cycle pulse at $\omega=0.1\omega_{0}$ (1D model): The function $j_{\omega}(x)j^{*}_{\omega}(y)$ is plotted, which determines the coherent photon emission probability via Eq. (16). Blue line indicates the tunnel exit $x_{e}=\kappa^{2}/(2E_{0})$. The region inside the blue sector shows nonnegligible contribution to the radiation.
Figure 4: Spontaneous radiation spectrum during tunneling ionization: (a) asymmetric few-cycle laser pulse, (b) symmetric laser pulse, in both case $N=2$.
We describe the electron state $\Psi$ in Eq. (5) during tunneling ionization via the strong-field approximation (SFA) [25]:
$\Psi=\psi^{(0)}+\psi^{(1)}+\psi^{(2)},$ (6)
@@
where the zeroth-order term $\psi^{(0)}$ coincides with the unperturbed bound state $\psi^{(a)}$ in the atomic potential. In the first-order SFA
$\psi^{(1)}=i\int^{s}dt\int d\mathbf{p}\psi^{f}_{\mathbf{p}}\langle\psi^{f}_{\mathbf{p}}(t)|\mathbf{r}\cdot\mathbf{E}(t)|\psi^{a}(t)\rangle$ (7)
@@
with the unperturbed continuum state $\psi^{f}$ in the laser field, Volkov state [26], and in the second-order SFA [25]:
$\psi^{(2)}$ $=$ $i\int^{s}dr\int^{r}dt\int d\mathbf{p}\int d\mathbf{q}\psi^{f}_{\mathbf{p}}$ (8)
$\times\langle\psi^{f}_{\mathbf{p}}(r)|V|\psi^{f}_{\mathbf{q}}(r)\rangle\langle\psi^{f}_{\mathbf{q}}(t)|\mathbf{r}\cdot\mathbf{E}(t)|\psi^{a}(t)\rangle.$
@@
The wave function $\psi^{(1)}$ describes the electron dynamics in the continuum, including the under-the-barrier dynamics, and the motion in the laser field after leaving the barrier. In Eq. (5) we can identify different probability currents $j(s)=j_{0,1}(s)+j_{1,0}(s)+j_{1,1}(s)$, where
$j_{0,1}(s)$ $=$ $\langle\psi^{(0)}(s)\left|\hat{p}_{x}\right|\psi^{(1)}(s)\rangle,$
$j_{1,0}(s)$ $=$ $\langle\psi^{(1)}(s)\left|\hat{p}_{x}\right|\psi^{(0)}(s)\rangle,$ (9)
$j_{1,1}(s)$ $=$ $\langle\psi^{(1)}(s)\left|\hat{p}_{x}\right|\psi^{(1)}(s)\rangle.$
The second-order terms $j_{0,2}$, $j_{2,0}$ describe the contribution to the radiation due to the under-the-barrier recollisions, which are usually small in the net radiation, see Sec. III. The radiation via the combination of $j_{01}$ and $j_{10}$ emerges due to the recombination to the bound state and, therefore, describes three-step HHG, including below-threshold harmonics. The radiation via $j_{11}$ term includes the effect of the electron transition to the continuum and the continuum dynamics, which will be the main point of our study.
For the physical interpretation, we introduce an heuristic model similar to Babushkin [10], to simulate the radiation only due to the continuum dynamics excluding the under-the-barrier contribution. In this model we evaluate the ionized electron probability current in the continuum, consisting of the tunnel ionized electron density via Amosov-Delone-Krainov (ADK) theory [27, 28] multiplied by the laser driven velocity:
@@
$\mathbf{j}_{B}(s)=\int^{s}dt[\mathbf{A}(s)-\mathbf{A}(t)]\,\frac{\kappa^{2}}{2}\frac{E_{0}}{E_{a}}\exp\left[-\frac{2E_{a}}{3|\mathbf{E}(t)|}\right],$ (10)
with the atomic field $E_{a}=\kappa^{3}$, $\kappa=\sqrt{2I_{p}}$, and the atomic ionization energy $I_{p}$, and calculate the corresponding spectrum via Eq. (5).

### v2

# Role of quantum dynamics in coherent and incoherent radiation during tunneling ionization

@@ Abstract
Radiation during strong-field tunneling ionization is investigated. The spontaneous as well as the coherent components of the radiation are calculated describing the ionization dynamics via the strong field approximation and the role of the quantum dynamics at tunneling is analyzed. The competition between different mechanisms in different spectral regions is examined. Transition-like radiation (Brunel radiation) is dominated at low-frequencies, Thomson scattering at the laser frequency, and radiative recombination via the three-step process at high-order harmonics. To distinguish the role of the quantum dynamics, simple man Drude models are developed for the coherent as well as for spontaneous radiation, which are based on the electron trajectory out of the tunneling barrier. The quantum dynamics is shown to modify the coherent Brunel radiation for near-zero-frequencies in asymmetric laser pulses. The significant role of free-free transitions is demonstrated for the spontaneous radiation in the low-frequency region.
@@ I Introduction
It is well known that tunneling ionization is accompanied by high-order harmonic generation (HHG) via the three-step model [1], when the recombination of the ionized electron wave packet to the bound state at the recollision after the excursion in the laser field and gaining energy results in the emission of a high energy photon [2, 3, 4]. At the recollision, radiation may emerge also due to electron scattering, rather than recombination, yielding free-free HHG [5]. In the macroscopic picture, the ionization of a gas target results in the creation of an electron current after the interaction with a laser pulse, which can be a source of low frequency THz radiation, with the emitted frequency determined by the interaction time [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The same mechanism can work also in the case of a single atom producing THz radiation in asymmetric laser pulses, e.g. in a two-color laser field [20, 21]. The origin of this radiation is similar to the transition radiation, when the electron velocity [22] (or dielectric constant of a medium where the electron propagates [23]) abruptly changes. A similar effect exists in plasma physics – Brunel radiation [24, 25], when an electron in the plasma is dragged out of the plasma. In the case of tunneling ionization, the electron is initially in a stationary bound state, while abruptly starting to move after tunneling.
In the tunneling regime of strong-field ionization, the bound electron tunnels through the barrier created by the laser and the Coulomb field of the atomic core. However, this common simple dynamics can continue to evolve in a nontrivial way. The electron traversing the classically forbidden region may reflect from the potential barrier surface, propagate back and scatter off the core, and then finally tunnel out of the barrier. Although this nontrivial path has no large probability, it has consequences, inducing the tunneling time delay when interfering with the direct ionization path [26, 27], and it can be responsible for driving Freeman resonances in the strong field regime [28], which has recently been confirmed experimentally [29]. The question arises whether the signatures of the quantum dynamics at tunneling, in particular, those of the under-the-barrier recollisions, can be observed in the radiation accompanying tunneling ionization.
@@
In contrast to the coherent radiation discussed above, emission of spontaneous low-order harmonic radiation during the bound-continuum transition with the origin similar to Brunel radiation has been recently shown in [30], with the typical frequency determined by the Keldysh tunneling time [31]. Finally, harmonic radiation at tunneling ionization arises also from Thomson scattering of the laser wave off the ionized free electron wave packet, with higher harmonics than the fundamental one suppressed in the nonrelativistic regime.
@@
Coherent radiation should be distinguished from incoherent radiation of the same origin. For instance, the spontaneous recombination of a high-energy electron to an atomic bound state results in a photon emission of a random phase, which is an incoherent process [32]. However, when the recombining electron is released from the bound state and recombines into the same bound state driven by a laser field, the HHG is coherent with the phase determined by the coherent superposition of the ground state wave function with the ionization continuum wave packet. Similarly, the coherent part of radiation due to Compton scattering of a strong laser wave from an ionized electron wave packet [33, 34] is phase-matched in forward direction, while the incoherent radiation has a broad angular distribution [35, 36, 37, 38, 39]. For the incoherent process, in contrast to the coherent one, the final state of the electron does not exist initially. The coherent radiation intensity scales as $\sim{\cal N}^{2}$ with the number of emitters $\cal N$}, due to the phase matching of the radiation, while for the incoherent radiation interference is suppressed and the total intensity scaling is $\sim{\cal N}$ [40]. The coherent radiation can be calculated using the quantum mechanical expectation value of the current density within the classical radiation theory [35].
@@
In this paper we investigate the role of the under-the-barrier dynamics for the spontaneous and coherent Brunel radiation during tunneling ionization. The ionization quantum dynamics is described by strong field approximation (SFA) which implicitly includes the under-the-barrier path. Additionally, we develop simple man Drude models for the coherent as well as for spontaneous radiation, which include only the electron trajectory out of the tunneling barrier. Comparison of the results of the Drude model with those of SFA allows us to single out the contribution of the under-the-barrier path in the radiation spectra. We analyze how the different mechanisms of the radiation during the strong field ionization are scaled in different spectral regions of emission and the feasibility of their observation in a coherent or incoherent form.
@@
The structure of the paper is the following. In Sec. II the coherent radiation during tunneling ionization is calculated using the SFA. Spontaneous emission at the tunneling ionization is discussed in Sec. III. Relation of our calculations to an experiment is considered in Sec. IV. We give an order of magnitude estimation of radiation intensities in Sec. V. Our conclusions are given in Sec. VI.
@@ II Coherent radiation during tunneling ionization
II.1 Radiation via SFA current
@@
We calculate coherent radiation emitted during tunneling ionization of a laser-driven hydrogenlike atom in the nonrelativistic regime. The dynamics of the tunneling ionization is treated by the SFA, which provides the current density of the ionizing electron. The latter is employed in Maxwell equations to calculate the intensity of coherent radiation.
The strong-field ionization dynamics is described by the Schrödinger equation
$i\partial_{t}|\Psi(t)\rangle=\hat{H}\,|\Psi(t)\rangle,$ (1)
with the Hamiltonian
$\hat{H}=\hat{\mathbf{p}}^{2}/2+\hat{V}(r)+\mathbf{r}\cdot\mathbf{E}(t)$ (2)
where $\hat{V}(\mathbf{r})=2\pi/\kappa\delta(\mathbf{r})(1+\partial_{r})$ is the regularized atomic zero-range potential, and $\mathbf{E}(t)$ the laser electric field. The interaction with the laser field is described in the length gauge which is the most suitable physically in the SFA [41, 42]. Atomic units are used throughout.
We seek the solution for the electron wave function $\Psi$ via the iterative SFA expansion [43]:
$|\Psi(t)\rangle=|\psi^{(0)}(t)\rangle+|\psi^{(1)}(t)\rangle+|\psi^{(2)}(t)\rangle,$ (3)
@@
where the zeroth-order term $|\psi^{(0)}\rangle$ coincides with the unperturbed bound state $|\psi^{(a)}\rangle$ in the atomic potential. The first-order SFA wave function describes the direct tunneling path:
$|\psi^{(1)}(t)\rangle=-i\int^{t}dt^{\prime}\int d^{3}\mathbf{p}\,|\psi^{(f)}_{\mathbf{p}}(t)\rangle\langle\psi^{(f)}_{\mathbf{p}}(t^{\prime})|\mathbf{r}\cdot\mathbf{E}(t^{\prime})|\psi^{(a)}(t^{\prime})\rangle$
@@
with the continuum Volkov state $|\psi^{(f)}(t)\rangle$ in the laser field [44], and in the second-order SFA [43]:
$|\psi^{(2)}(t)\rangle$ $=$ $-\int^{t}dt^{\prime\prime}\int^{t^{\prime}}dt^{\prime}\int d^{3}\mathbf{p}\int d^{3}\mathbf{q}\,|\psi^{(f)}_{\mathbf{p}}(t)\rangle$ (5)
$\times\langle\psi^{(f)}_{\mathbf{p}}(t^{\prime\prime})|V|\psi^{(f)}_{\mathbf{q}}(t^{\prime\prime})\rangle\langle\psi^{(f)}_{\mathbf{q}}(t^{\prime})|\mathbf{r}\cdot\mathbf{E}(t^{\prime})|\psi^{(a)}(t^{\prime})\rangle.$
@@
The states $|\psi^{(1)}\rangle$ and $|\psi^{(2)}\rangle$ describe the electron dynamics in the continuum and at recollisions, including the under-the-barrier dynamics, and the motion in the laser field after leaving the barrier.
From previous studies it is known [35] that the coherent radiation can be calculated interpreting the quantum current density as a classical one within the classical radiation theory. We calculate the quantum current density for the tunneling electron $\mathbf{j}(\mathbf{r},t)=\frac{1}{2}\{\Psi(\mathbf{r},t)^{*}\hat{\mathbf{p}}\Psi(\mathbf{r},t)+\rm c.c.\}$, interpret it as a classical current and use it to calculate the emitted coherent radiation [40] within the standard gauge invariant representation of Lienard-Wiechert radiation theory. From the latter the photon emission probability reads [45]:
@@
$dw=\frac{\alpha}{(2\pi)^{2}}\frac{d^{3}\mathbf{k}}{k^{3}}\left|\int dt\int\,d^{3}\mathbf{r}\,\mathbf{k}\times\mathbf{j}(\mathbf{r},t)e^{i\omega t-i\mathbf{k}\cdot\mathbf{r}}\right|^{2},$ (6)
where $\alpha$ is the fine structure constant, $\omega$ and $\mathbf{k}$ are the emitted photon energy and momentum, respectively, and $k=|\mathbf{k}|$. In the nonrelativistic regime the dipole approximation for the emission is relevant, $\mathbf{k}\cdot\mathbf{r}\ll 1$. Then, the angular integration of the emitted photon probability can be carried out yielding:
$\frac{dw}{d\omega}=\frac{2\alpha\omega}{3\pi c^{2}}\left|\int dt\,\mathbf{J}(t)\exp(i\omega t)\right|^{2},$ (7)
where $\mathbf{J}(t)=\int d^{3}\mathbf{r}\,\mathbf{j}(\mathbf{r},t)=\langle\Psi(t)\left|\hat{\mathbf{p}}\right|\Psi(t)\rangle$ is the expectation value of the electron probability current, and $c$ is the speed of light. We can identify different probability currents $\mathbf{J}(t)=\mathbf{J}_{01}(t)+\mathbf{J}_{10}(t)+\mathbf{J}_{11}(t)$, where
$\mathbf{J}_{01}(t)$ $=$ $\langle\psi^{(0)}(t)\left|\hat{\mathbf{p}}\right|\psi^{(1)}(t)\rangle,$
$\mathbf{J}_{10}(t)$ $=$ $\langle\psi^{(1)}(t)\left|\hat{\mathbf{p}}\right|\psi^{(0)}(t)\rangle,$ (8)
$\mathbf{J}_{11}(t)$ $=$ $\langle\psi^{(1)}(t)\left|\hat{\mathbf{p}}\right|\psi^{(1)}(t)\rangle.$
The radiation via the combination of $\mathbf{J}_{01}$ and $\mathbf{J}_{10}$ emerges due to the recombination to the bound state and, therefore, describes three-step HHG, including below-threshold harmonics. The radiation via $\mathbf{J}_{11}$ term includes the effect of the electron transition to the continuum and the continuum dynamics, which will be the main point of our study. The second-order terms $\mathbf{J}_{02}$, $\mathbf{J}_{20}$ describe the contribution to the radiation due to the under-the-barrier recollisions, which are usually small in the net radiation, and the observation of its signatures requires more subtle analysis, see Sec. IIIC.
The expression for the photon emission probability of Eq. (7) can be equivalently derived via the quantum perturbation theory, treating the interaction of the electron with the photon field in the velocity gauge $H_{\rm rad}=\hat{\mathbf{A}}(t)\cdot\hat{\mathbf{p}}$ by perturbation:
$dw$ $=$ $\left|m_{\mathbf{k}}\right|^{2}d^{3}\mathbf{k}$ (9)
$m_{\mathbf{k}}$ $=$ $-i\int dt\langle\Psi(t),1_{\mathbf{k}}|\hat{\mathbf{A}}(t)\cdot\hat{\mathbf{p}}|\Psi(t),0_{\mathbf{k}}\rangle,$
where $\hat{\mathbf{A}}=-\hat{\mathbf{E}}(t)/\omega$ is the vector potential operator for the emitted photon, and $\hat{\mathbf{E}}(\mathbf{r},t)=\sum_{\mathbf{k}}\sqrt{\omega}/(2\pi)\hat{\mathbf{e}}\left(a_{\mathbf{k}}e^{i\mathbf{k}\cdot\mathbf{r}-i\omega t}+a^{+}_{\mathbf{k}}e^{-i\mathbf{k}\cdot\mathbf{r}+i\omega t}\right)$ is the field of the emitted photons in second quantization, with the creation and annihilation operators $a^{+}_{\mathbf{k}}$ and $a_{\mathbf{k}}$. As the classical radiation theory is gauge invariant, we take its result as a benchmark to choose the gauge of the quantum perturbation theory with respect to the radiation Hamiltonian. Thus, we use different gauges in description of different interactions: the length gauge to describe the electron interaction with the laser field, justified by the SFA experience, and the velocity gauge for the perturbative treatment of the interaction with the emitted photon field, justified by the matching with the Lienard-Wiechert theory.
II.2 Simple man Drude model for Brunel radiation
The radiation via the current $\mathbf{J}_{11}$ in Eq. (7) describes radiation during tunneling ionization due to free-free transitions. This includes, first of all, Brunel radiation, which arises due to the bound electron abruptly starting to move after tunneling, and secondly, Thomson scattering of the laser wave on the ionized free electron, which contributes mostly at the laser frequency. The Bremsstrahlung in the applied short-range atomic potential is suppressed. We are mostly interested in Brunel radiation. We will see that the current $\mathbf{J}_{11}$ contains contributions from the continuum dynamics after the tunneling, as well as from the under-the-barrier path. To clearly single out the effect of the under-the-barrier dynamics for the Brunel radiation, we introduce a simple man model dubbed as Drude model in Ref. [20], and simulate the radiation only due to the continuum dynamics excluding the under-the-barrier contribution.
In this model we evaluate the ionized electron probability current in the continuum $\mathbf{J}_{\rm D}(t)$ at the time instant $t$, consisting of the tunnel ionized electron probability contribution at previous times $t^{\prime}<t$, multiplied by the laser driven velocity after the tunneling. We use the ionization probability via the Perelomov-Popov-Terent’ev (PPT) theory [46, 47], including nonadiabatic corrections:
$W(t^{\prime})=\frac{\kappa^{2}}{2}\frac{|\mathbf{E}(t^{\prime})|}{E_{a}}\exp\left[-\frac{2E_{a}\left(1-\frac{\gamma^{2}}{10}\right)}{3|\mathbf{E}(t^{\prime})|}\right],$ (10)
where $E_{a}=\kappa^{3}$ is the atomic field, $I_{p}$ the atomic ionization energy, $\kappa=\sqrt{2I_{p}}$, $\gamma=\kappa\omega_{0}/E_{0}$ Keldysh parameter. The laser driven velocity after the tunneling is
$\mathbf{v}(t,t^{\prime})=\mathbf{A}(t)-\mathbf{A}(t^{\prime}).$ (11)
Thus, the ionized electron probability current in the continuum reads:
$\mathbf{J}_{\rm D}(t)=\int^{t}dt^{\prime}\,\mathbf{v}(t,t^{\prime})W(t^{\prime}).$ (12)
Using $\mathbf{J}_{\rm D}(t)$, the radiation spectra with the Drude model is calculated using classical radiation theory via Eq. (7).
Within this model, the radiation has a spike in the near-zero-frequency domain in the case of an asymmetric pulse, which has a simple explanation. Initially, the electron is at rest at the core, then it tunnels out during the current formation time $\tau$, and is accelerated in the continuum due to the laser field up to its final velocity $v$. The current associated with this motion can be parameterized as
$J(t)\sim v\,\,[1+\text{erf}(t/\tau)],$ (13)
with the Fourier component of the current $J_{\omega}\sim v\exp(-\omega^{2}\tau^{2}/4)/\omega$. Thus, from Eq. (7)
$\frac{dw}{d\omega}\sim\frac{v^{2}}{\omega}e^{-\omega^{2}\tau^{2}/2},$ (14)
and the spectrum has a singularity at zero-frequency. Note that the Drude model exaggerates the singularity because it does not account accurately for the formation time $\tau$ of the current, because the error-function of Eq. (13) is replaced by a Heaviside-theta function.
Figure 1: Coherent radiation spectra during tunneling ionization with $E_{0}=0.1$, $\omega_{0}=0.05$, and $N=12$: (a) in the asymmetric laser pulse ${\cal A}=0.1$ [Eq. (15)], (b) in the symmetric laser pulse ${\cal A}=0$. Comparison of the Brunel radiation via $J_{11}$ with the three-step HHG ($J_{01}+J_{10}$) and with the Drude model ($J_{\rm D}$). The total radiation with the full current $J=J_{01}+J_{10}+J_{11}$ is given for comparison.
Figure 2: Coherent radiation spectra during tunneling ionization in the asymmetric laser pulse: (first column) $E_{0}=0.1$ [zoomed-in view Fig. 1(a)]; (second column) $E_{0}=0.075$. Comparison of the Brunel radiation via $J_{11}$ with the three-step HHG ($J_{01}+J_{10}$) and with the Drude model ($J_{\rm D}$).
II.3 Coherent Brunel radiation
II.3.1 Calculation of the current